# AL-Bench: A Benchmark for Automatic Logging

### v2

@@ Abstract.
Logging, the practice of inserting log statements into source code, is critical for improving software reliability. Recently, language model-based techniques have been developed to automate log statement generation based on input code. These tools show promising results in their own evaluation. However, current evaluation practices in log statement generation face significant challenges. The lack of a unified, large-scale dataset forces studies to rely on ad-hoc data, hindering consistency and reproducibility. Additionally, assessments based solely on metrics like code similarity fail to reflect real-world effectiveness. These limitations underscore the need for a comprehensive public benchmark to standardize evaluation.
This paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools. AL-Bench includes a high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements and introduces a novel dynamic evaluation approach. Different from the existing evaluations that focus only on components of log statements like code similarity, AL-Bench assesses both the compilability of the code with inserted log statements and the effectiveness of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice. AL-Bench reveals significant limitations in the state-of-the-art tools. The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases. In addition, even the best-performing tool only achieves 0.213 cosine similarity between the runtime logs produced by the generated log statements and the ground-truth log statements. The results reveal substantial opportunities to further enhance the development of automatic logging tools.
Logging, Software Maintenance, Logging Benchmark
@@
As software grows in size and complexity, logging has become increasingly essential to ensuring software reliability (He et al., 2021; Chen and Jiang, 2021). Logging means writing log statements into the source code, which generate runtime logs that record valuable information for a range of downstream tasks such as anomaly detection (Zhang et al., 2019; Lou et al., 2010; He et al., 2016; Du et al., 2017; Zhang et al., 2019), fault diagnosis (Zou et al., 2016), root cause analysis (Amar and Rigby, 2019; He et al., 2018b; Lin et al., 2016), and program verification (Ding et al., 2015; Shang et al., 2013). The effectiveness of these downstream tasks heavily relies on the quality of the software logs (He et al., 2018a). Therefore, appropriate logging is essential to capture critical behaviors during software operation (Yuan et al., [n. d.]).
@@
Numerous logging tools have been developed to assist software developers by automatically suggesting log statements based on provided code snippets (Mastropaolo et al., 2022; Liu et al., 2021; Xu et al., 2024; Xie et al., 2023). As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded. Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE (Mastropaolo et al., 2022) is the first end-to-end tool to integrate positioning, level selection, and message generation. UniLog (Xu et al., 2024) employed a warm-up and in-context learning (ICL) strategy to enhance performance. FastLog (Xie et al., 2023) improved the generation efficiency while maintaining precision. LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance. SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet. These studies typically evaluate performance using ad-hoc data splitting from the entire dataset, focusing on metrics such as the accuracy of log statement components, including position, level, and message. Machine translation metrics, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), are also employed to evaluate the quality of generated log messages. Although these evaluations offer valuable insights into logging tool performance, the use of low-quality data and incomplete assessments undermines the reliability of the results.
First, loose standards collection and inappropriate clean strategy compromise the quality of evaluation data and effectiveness in assessing the performance. Previous evaluation datasets were typically created by splitting the entire dataset. The data selection rules of the entire dataset are commonly loose standards to ensure sufficient data for training. These criteria fail to ensure the quality and consistency of the data. Moreover, to accommodate the limitations of the tools, they filtered out all instances exceeding 512 tokens, ignoring the long code snippets that are commonly encountered in real-world development environments. This approach undermines the overall effectiveness and real-world applicability of the evaluation results, as it does not accurately reflect the true complexity of software projects.
Second, the current evaluation method does not verify whether the generated log statements are compilable. Generating the compilable log statements is the first requirement when applying automatic logging tools in practice. To relieve developers from the heavy effort required to design and maintain log statements (Chen and Jiang, 2019, 2017), the basic requirement is to ensure that our tool can be seamlessly integrated into the DevOps process without introducing additional errors that require extra debugging effort from developers. Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or use undefined variables. Evaluating the compilability of predicted log statements reflects the effectiveness and reliability of logging tools, which are essential for practical use.
Third, the evaluation method cannot evaluate the quality of runtime logs generated by the predicted log statements. Current evaluation methods assess the performance of tools based on the correctness of individual components of log statements. However, current metrics struggle to accurately reflect the quality of runtime logs in the real execution environment. Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-line shift from the ground truth or a mismatch in verbosity level. For instance, a minor difference in verbosity levels (e.g., debug vs. info) can cause critical information missed due to log level threshold settings in the source code. Therefore, we need to evaluate the quality of log statements in a real execution environment, with the goal of obtaining appropriate logs in specific scenarios, rather than merely focusing on the correctness of individual components or relying on statistical metrics to reflect the performance of logging tools.
Our Work. To address these challenges, we introduce AL-Bench: a comprehensive benchmark featuring a large-scale, diverse dataset and a novel approach for evaluating both the compilability and runtime logs produced by generated log statements. Our dataset comprises 42,224 instances collected from 10 popular, high-quality GitHub projects (GitHub, [n. d.]) spanning different domains with varying logging requirements, providing a robust foundation for evaluating automatic logging methods. Beyond the dataset, AL-Bench introduces a novel evaluation method called dynamic evaluation, which involves reintegrating generated log statements into real project code, followed by recompiling and executing them. This process allows for a realistic assessment of both compilability and runtime logs, highlighting major limitations in state-of-the-art tools: even the best tools fail to compile in 20.1% of cases, and the logs produced by generated log statements show only 0.213 cosine similarity to ground-truth logs. Through its rigorous and standardized evaluation approach, AL-Bench bridges the gap between real-world logging requirements and prior assessments, highlighting substantial opportunities to further advance the development of automatic logging tools.
This paper’s contributions are summarized as follows:
@@
    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub (GitHub, [n. d.]) projects, spanning various domains with differing logging requirements.
@@
    We propose a novel dynamic evaluation approach that assesses generated log statements in real-world settings by reintegrating them into projects to evaluate their compilability and resulting runtime logs.
@@
2. background and motivation
This section introduces the overview of the current research on automatic logging tools, followed by the shortcomings in the evaluation work, and explains the motivation for AL-Bench.
2.1. Log Statement Generation
@@
Recently, Mastropaolo et al. (Mastropaolo et al., 2022) proposed the first end-to-end tool LANCE to generate complete log statements based on T5 (Raffel et al., 2019). Following LANCE, Xu et al. (Xu et al., 2024) proposed UniLog to adapt ICL and warm-up strategy to enhance the LLM ability for generating log statements. Xie et al. (Xie et al., 2023) proposed FastLog increase the generation time while keeping the accuracy, and Mastropaolo et al. (Mastropaolo et al., 2023) further proposed LEONID with a combination of Deep Learning (DL) and Information Retrieval (IR) achieving a better performance.
While end-to-end automatic logging tools have demonstrated promising results in their respective evaluations, our analysis reveals notable issues in both the datasets used and the evaluation methods employed.
Figure 2. The low-quality examples in the LANCE dataset follow the strictest data collection rules.
2.2. Limitations of Evaluation Methodology
2.2.1. Evaluation Data
Current evaluation datasets are ad-hoc, derived from splitting the entire collected data according to loose standards. More specifically, to ensure a sufficient amount of training data, previous studies had to adopt lenient rules for data collection. Additionally, some tools, to accommodate the limitations of their backbone models, further filter out data exceeding 512 tokens in length. These strategies compromise data quality, reducing the reliability and effectiveness of evaluation results. For example, LANCE adopted the strictest data collection rules among all tools, requiring a minimum of 500 commits, 10 contributors, and 10 stars, while also excluding forked repositories on GitHub (GitHub, [n. d.]). However, even in the LANCE dataset, low-quality data is pervasive. As shown in Figure 2, several patterns of low-quality logging practices persist. These patterns include the duplication of variables within a single log statement, creating unnecessary redundancy; the inclusion of empty strings or meaningless content, resulting in uninformative messages; excessive use of special characters or punctuation, making printed logs difficult to parse; and mismatched logging levels, where critical messages are assigned inappropriate severities, leading to misclassification (Adzic, 2006; Chen and Jiang, 2017). Using low-quality data as an evaluation dataset fails to produce reliable results. Such results do not align with the expectation that logging tools should generate high-quality log statements within the code. Furthermore, to accommodate the input length limitations of these tools, previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023) have filtered out instances longer than 512 tokens. While this data-cleaning strategy reduces the complexity of evaluation data, it also introduces significant biases by excluding longer code snippets, which are common in real-world development environments. This approach simplifies the evaluation process but fails to capture the full scope of challenges that logging tools would face when dealing with complex and extended code bases. Consequently, the evaluation results may not accurately reflect the tools’ ability to handle the demands of real-world software development, where longer snippets and more intricate code structures are prevalent. Therefore, there is a pressing need for a public, large-scale, high-quality, and diverse benchmark dataset that can better represent real-world codebases and provide a standardized platform for evaluating automatic logging tools.
Figure 3. An example of the limitation of static evaluation, where only the quality of log statements is considered. It cannot assess the compilability of the code or quality of the printed runtime logs.
2.2.2. Evaluation Method
The current evaluation method focuses on assessing model performance by comparing the accuracy of each log statement component (i.e., position, verbosity level, message). We refer to this as static evaluation in this paper. Figure 3 provides an example of static evaluation to demonstrate how it works. In this example, only the verbosity level matches, so the matched level count will increase by one, and BLEU scores will be used to calculate the average score. Finally, the matched portion of each component and the average BLEU score will be combined to reflect the performance of the tools. These statistical metrics provide insights into evaluating the quality of predicted log statements; however, this approach still does not align with the goal of ensuring that the predicted log statements generate high-quality logs.
Figure 4. An example of compilation failure: Unreachable Statements
First, it cannot assess the compilability of predicted log statements. Compilability is a fundamental requirement for generating appropriate logs. This capability ensures that log statements can be applied to the source code without introducing compilation errors, avoiding wasting developers’ efforts to debug for the predicted log statements. However, the current evaluation methodology, which focuses solely on comparing the individual components of log statements, cannot identify predicted log statements that use undefined variables or contain incorrect code syntax which could introduce compilation errors. As shown in Figure 4, the predicted log statement is injected into an unreachable position. Applying static evaluation would fail to recognize that this prediction could cause compilation errors. Instead of classifying it as a completely incorrect prediction, the evaluation would count it as a level match instance and assign it a BLEU score of 55.03, contributing to the statistical value. This could introduce significant bias when evaluating the performance, appearing favorable in metrics but performing poorly in actual usage.
Second, it cannot assess the runtime logs generated by predicted log statements. The goal of writing log statements is to obtain the appropriate logs during software execution so that the quality of log statement is ultimately decided by the quality of runtime logs. Although comparing the correctness of log statement components provides insights into assessing the quality of logs, it can introduce bias when using statistical metrics that focus solely on log statements to evaluate the performance of logging tools. A simple example is a mismatched verbosity level between the predicted log statement and the ground truth. In a case where only the verbosity level is mismatched while the other components are exactly matched, this instance would receive a high score in static evaluation. However, in a real-world scenario, for important error logs, if the predicted level is below the threshold, the logs will not be generated, regardless of how critical the message is. Conversely, for debug logs, if the predicted level exceeds the threshold, these logs could be printed in the production environment, leading to issues such as sensitive information leakage and increased storage costs. These two situations demonstrate that even a shift in verbosity level can turn a log statement into a completely bad case. A more complicated case is presented in Figure 5. Under static evaluation, this prediction with only a shift in two lines in position, also could be rewarded highly in statistic metrics. However, it is completely incorrect: placing the log statement at line 2 risks a NullPointerException if the timer object is null when timer.get() is called. This would cause the program to crash during execution, meaning that not only would the logs fail to be generated, but the entire program could terminate unexpectedly. In real-world scenarios, this kind of error is critical and goes unnoticed in static evaluations. While static evaluation rewards this prediction for structural correctness, it fails to account for the actual behavior of the code when executed. Therefore, this example reinforces the need for more comprehensive evaluation methods that go beyond code-level metrics and consider the real-world implications of log statements.
In conclusion, a log statement’s quality should be determined by its ability to avoid introducing extra errors and generate meaningful, contextually appropriate logs.
Figure 5. An example of the limitation of static evaluation. The prediction differs from the ground truth by a slight upward shift of two lines. It causes a NullPointerException when timer is null and fails to generate logs.
Insights: The motivating study underscores major limitations in current evaluation efforts, including the lack of a diverse, high-quality dataset and methods that do not meet real-world needs. To improve log statement evaluation, we should collect a high-quality, diverse dataset and use an execution-based method to directly evaluate the runtime logs generated by log statements.

### v3

@@ Abstract.
Logging, the practice of inserting log statements into source code, is critical for improving software reliability. Recently, language model-based techniques have been developed to automate log statement generation based on input code. While these tools show promising results in prior studies, the fairness of their results comparisons is not guaranteed due to the use of ad hoc datasets. In addition, existing evaluation approaches exclusively dependent on code similarity metrics fail to capture the impact of code diff on runtime logging behavior, as minor code modifications can induce program uncompilable and substantial discrepancies in log output semantics. To enhance the consistency and reproducibility of logging evaluation, we introduce AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools. AL-Bench includes a large-scale, high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements. Moreover, it introduces a novel dynamic evaluation methodology to provide a run-time perspective of logging quality in addition to the traditional static evaluation at source code level. Specifically, AL-Bench not only evaluates the similarity between the oracle and predicted log statements in source code, but also evaluates the difference between the log files printed by both log statements during runtime. AL-Bench reveals significant limitations in existing static evaluation, as all logging tools show average accuracy drops of 37.49%, 23.43%, and 15.80% in predicting log position, level, and message compared to their reported results. Furthermore, with dynamic evaluation, AL-Bench reveals that 20.1%-83.6% of these generated log statements are unable to compile. Moreover, the best-performing tool achieves only 21.32% cosine similarity between the log files of the oracle and generated log statements. These results underscore substantial opportunities to advance the development of automatic logging tools. We believe this work establishes a foundational step in furthering this research direction.
Software Maintenance, Logging, Benchmark
@@
As software grows in size and complexity, logging has become increasingly essential to ensuring software reliability (He et al., 2021; Chen and Jiang, 2021). Logging means writing log statements into the source code, which generate runtime logs that record valuable information for a range of downstream tasks such as anomaly detection (Zhang et al., 2019; Lou et al., 2010; He et al., 2016; Du et al., 2017; Zhang et al., 2019), fault diagnosis (Zou et al., 2016; Xu et al., 2023, 2024c), root cause analysis (Amar and Rigby, 2019; He et al., 2018b; Lin et al., 2016; Xu et al., 2025), and program analysis (Ding et al., 2015; Shang et al., 2013; Xu et al., 2024b). The effectiveness of these downstream tasks heavily relies on the quality of the software logs (He et al., 2018a). Therefore, appropriate logging is essential to capture critical behaviors during software operation (Yuan et al., [n. d.]).
@@
To help software developers implement effective and efficient logging practices, researchers have been focusing on exploring methodologies for automatic logging in recent years. As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded. To address the automatic logging tasks, numerous logging methods have been proposed thanks to the advanced development of the Large Language Model (LLM). For example, LANCE (Mastropaolo et al., 2022) was introduced as the first end-to-end logging method, seamlessly integrating the recommendation of log statement position, verbosity level, and message content. Building on this foundation, UniLog (Xu et al., 2024a) employed a warm-up and in-context learning (ICL) strategy to enhance performance. FastLog (Xie et al., 2024) improved the generation efficiency while maintaining precision. LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance. SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet.
However, current logging research faces a significant challenge: The absence of a standardized evaluation dataset and methodology hinders fair comparisons between different logging methods. This issue manifests in two key aspects: (1) The test sets of different logging methods are often derived from their own training data, leading to inconsistent experimental performance and hindering fair comparisons. (2) The evaluation metrics used for different logging methods are not fully consistent, resulting in varying advantages across different dimensions. Given these challenges, we argue that establishing a comprehensive logging benchmark is essential. However, it is infeasible to naively construct a benchmark via ensembling existing evaluation data and metrics due to two major limitations of existing evaluation datasets and methods:
First, existing evaluation datasets lack the quality required for reliable assessment. For example, 18.45% cases (2218/12020) in LANCE’s test set and 30.36% cases (2197/7237) in FastLog’s test set involve numerous bad patterns of log messages. As illustrated in Figure 2 these anomalies predominantly fall into the following types: (a) duplicated content, (b) empty string, (c) duplicated empty and special characters, (d) contents that mismatch the log levels, and (e) explicit type casting. These issues distort evaluation outcomes, as tools producing such flawed log statements tend to appear better on metrics despite having poor logging quality. Moreover, these datasets are restricted to cases with very short contents ($\leq$ 512 tokens), a limitation imposed by their use of foundation models (e.g., T5, PLBART), which support a maximum input of 512 tokens. This constraint not only reflects a compromise to accommodate their models but also inhibits the evaluation of realistic, function-level logging capabilities, as many industrial functions exceed this token limit. Consequently, the datasets fail to meet the requirements for assessing logging quality in real-world scenarios.
Second, current evaluation methods lack a fine-grained view for comprehensive assessment. Existing evaluations primarily focus on static code similarity between candidate log statements generated by logging methods and oracle log statements written by developers. However, the code compilability is entirely neglected and can not be measured by static evaluation. Our preliminary analysis revealed that 20%-80% of log statements generated by current methods fail to compile, a figure that has not been previously reported. Furthermore, such static similarity fails to capture logging performance during program execution. For instance, placing a log statement inside or outside a loop may differ by only one line in code but result in vastly different numbers of log entries at runtime. Conversely, moving a log statement above or below a comment may have no effect on the runtime logs despite a one-line difference in the code. Such runtime differences are overlooked in static evaluations. These limitations highlight fundamental flaws in the current evaluation approach.
Figure 2. Bad patterns in existing datasets (using instances from LANCE dataset as examples): (1) Duplicated Variable records the same information multiple times, costing redundant overhead for both printing and recording logs in runtime. (2) Empty String provides insufficient context, hindering the effectiveness of debugging through printed logs. (3) Unpredictable Character contains numerous meaningless special tokens, making both the logging methods hard to predict and their printed logs difficult to parse for downstream analysis (Adzic, 2006). (4) Wrong Verbosity Level often misleads developers for debugging and fault localization (Chen and Jiang, 2017). (5) Explicit Cast couple logs to variable type casting, might cause runtime type conversion errors and system crash (Chen and Jiang, 2017).
To address these two challenges, our key insights are: (1) Log statements from the most widely used and frequently updated repositories tend to be of higher quality with fewer bad logging patterns. (2) Log files printed in runtime can serve as a proxy for evaluating log statements quality with a more comprehensive view of program logging performance. Based on these insights, we introduce AL-Bench, the first unified logging benchmark involving a large-scale evaluation dataset and a comprehensive evaluation method for assessing both the static log statements and runtime log files printed via code execution. Our dataset comprises 21,804 instances mined from the 10 most popular and actively maintained GitHub projects (GitHub, [n. d.]), each with at least 10,000 stars, and 500 log-related issues spanning diverse domains and application scenarios. Furthermore, AL-Bench introduces dynamic evaluation in addition to traditional static evaluation for log statements. Specifically, it not only applies a suite of static metrics to quantify the quality of generated log statements but also reintegrates them into real project code, compiles, and executes them. This approach enables a realistic assessment beyond the traditional static approach, highlighting major limitations in existing logging methods: even the best-performing logging method fails to compile in 20.1% of cases, and the logs printed in runtime show only 21.32% cosine similarity to the oracle logs. By providing a unified, standardized, and comprehensive evaluation dataset and methodology, we believe AL-Bench establishes a foundational step toward advancing research in automatic logging.
To sum up, our contributions are shown as follows:
@@
    We collected a high-quality, diverse, and large-scale dataset comprising 21,804 instances from 10 popular, high-quality GitHub (GitHub, [n. d.]) projects, spanning various domains with differing logging requirements.
@@
    We propose a dynamic evaluation approach that reintegrates the generated log statements into projects to evaluate their printed runtime logs as a more fine-grained proxy for illustrating logging performance.
@@
2. Background
2.1. Related Work of Automatic Logging
@@
Recently, with the advanced capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), numerous LLM-based logging methods have been proposed. Specifically, LANCE (Mastropaolo et al., 2022) was introduced as the first end-to-end logging method, seamlessly integrating the recommendation of log statement position, verbosity level, and message content. Building on this foundation, UniLog (Xu et al., 2024a) employed a warm-up and in-context learning (ICL) strategy to enhance performance. FastLog (Xie et al., 2024) improved the generation efficiency while maintaining precision. LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance. SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet.
2.2. Philosophy of Logging Evaluation
Logging quality evaluation has always been a critical yet challenging problem in automatic logging (Bogatinovski et al., 2022; Li et al., 2019; Chen and Jiang, 2017, 2021; He et al., 2021). In an ideal scenario, high-quality logs should accurately capture key runtime state information of a software system while avoiding excessive logging that may lead to unnecessary consumption of computational and storage resources. Intuitively, the most direct approach to assessing logging quality is to evaluate how well the logs support specific downstream tasks (Yuan et al., [n. d.]). However, using downstream tasks as proxies to measure general-purpose logging quality in a scalable way is both theoretically and practically problematic for several reasons:
  1. (1)
    First, it is very challenging to cover all potential downstream tasks for logs exhaustively. There are countless log-based downstream tasks: beyond the tasks with clear task formulations (e.g., debugging, anomaly detection, root cause analysis), logs are also used for a wide range of tasks that lack a universal formulation (e.g., recording sequences of runtime events to monitor system status), often driven by the unique needs of each system’s developer (Zhang et al., 2019; Xu et al., 2024c; Amar and Rigby, 2019). Even if we only choose a fixed set of tasks with clear formulations to evaluate logging quality, these proxies can only assess task-specific aspects rather than provide a comprehensive measure of overall logging quality. This limitation makes them unreliable indicators of general-purpose logging quality.
  2. (2)
    Second, it is not practical to build testbeds for these downstream tasks across all software systems. While it might be theoretically possible to create a testbed for evaluating logging quality in common tasks like anomaly detection, assessing tasks like whether logs capture key information for event monitoring needs is challenging, as it depends heavily on subjective developer requirements. Moreover, even creating fault datasets to support log-based anomaly detection is also challenging since different software systems have unique fault characteristics, making scalable fault injection across diverse environments infeasible (Zhao et al., 2021).
  3. (3)
    Third, the trade-off between logging effectiveness and efficiency is another unmeasurable aspect of logging quality. Massive logging offers detailed runtime data for fault diagnosis but increases storage costs and resource usage, while sparse logging reduces expenses but risks missing critical information. System designers balance these tradeoffs primarily through operational experience from maintaining and debugging specific systems. However, such experiences are inherently context-dependent, varying across systems due to divergent functional, architectural, and operational priorities (Rong et al., 2023). Consequently, it remains challenging to establish a universal trade-off mechanism for assessing the appropriateness of logging strategies for each specific software.
As an alternative, logging research often adopts a similarity-based approach to evaluate logging quality. The core idea is that, since assessing log quality via downstream tasks is challenging, we can use high-quality log statements as references to evaluate new ones. Thus, by constructing benchmarks, we can check if a logging method generates log statements similar to these reference ones, making evaluation scalable with log statements from open-source software.
However, this similarity-based approach does have limitations. It may not accurately assess log statements of higher quality than the reference, as they could receive lower similarity scores simply for not matching closely enough. To mitigate this, reference log statements must be of sufficiently high quality, which poses a challenge in data collection. We argue that log statements from the most widely-used, time-tested projects are the best reference candidates. Additionally, current evaluations focus solely on static code, but small changes in code can lead to significant differences in log content. Therefore, evaluating both static code and the generated log files provides a more comprehensive measure of logging quality. These insights inform the development of our first logging benchmark: AL-Bench.